OpenStudy (anonymous):

I HAVE A FEW QUESTIONS. AND IF YOU HELPED ME THAT WOULD BE GREAT <333

OpenStudy (anonymous):

yeah tell me them

OpenStudy (anonymous):

Which is the direct linear variation equation for the relationship? y varies directly with x and y = 10 when x = 2. A. y = 5x B. y = x + 8 C. y = 2x + 6 D. y = x – 8 i think it's B

OpenStudy (anonymous):

your right

OpenStudy (anonymous):

thankyou <3 Which is the quadratic variation equation for the relationship? y varies directly with x2 and y = 48 when x = 2. A. y = 4x B. y = 12x2 C. y = 4x2 D. y = x2 + 25 i think its B

OpenStudy (anonymous):

um ... wait

OpenStudy (anonymous):

This is the equation for direct variation for quadratics. Plug in what we know: 48 = k(2)^2 What's 2^2?

OpenStudy (anonymous):

@onedirection_

OpenStudy (anonymous):

oh sorry @familiabarajas im back

OpenStudy (anonymous):

2^2 is 4

OpenStudy (anonymous):

theres your anwser

OpenStudy (anonymous):

so its y=4x

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

yeah no problem

OpenStudy (anonymous):

What is the range of Variable A for the cluster? ___ to ___ http://static.k12.com/calms_media/media/1549000_1549500/1549022/1/5b8019e8d400ba25fc4871a67b0856fe0c925aca/MS_IMC-141015-108155.jpg

OpenStudy (anonymous):

i think its 1 to 4 @familiabarajas (:

OpenStudy (anonymous):

i dont know this go with @dan815

OpenStudy (anonymous):

okay(: i just need someone to check it

OpenStudy (dan815):

range is the min height to max height

OpenStudy (dan815):

the lowest height is 1, how about the highest http://prntscr.com/6xazm7

OpenStudy (anonymous):

3

OpenStudy (anonymous):

oh and thankyou for helping me :3

OpenStudy (anonymous):

is it 1-3